OpenStudy (anonymous):

Suppose the demand for a product is given by D(p)=-4p+198. A) Calculate the elasticity of demand at a price of $21. (Give your answer to three decimal places.) Elasticity ? and what price do you have unit elasticity?

OpenStudy (anonymous):

21=-4p+198 4p=198-21 4p=177 p=44.25 is it correct if it will be the new price? or it should be q?